OpenStudy (anonymous):

In solving the equation (x + 1)(x – 2) = 4, Eric stated that the solution would be x + 1 = 4 => x = 3 or (x – 2) = 4 => x = 6 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

because the answers are -2 and 3

OpenStudy (anonymous):

the equation has to be set to = 0

OpenStudy (anonymous):

so you foil it out to get x^2-x - 2 -4

OpenStudy (anonymous):

becomes x^2-x -6 =0

OpenStudy (anonymous):

now factor it to get (x-3)(x+2)=

OpenStudy (anonymous):

0

OpenStudy (anonymous):

x = 3, -2

OpenStudy (anonymous):

1) we find x only when right side is zero.then we equalise factors to zero 2) but in this case we found x from x+1=4 ...now if we put this x in original equation then it can't satisfy. .because the x for which x+1 is equal to 4 ..for the same x ...your x-2 will be something different..so their product will never equal to 4 because its one part(x+1) itself gives 4.